please explain why

Suppose the Xavier values each apple at $1 and each banana at $2; Yolanda values each apple at $2 and each banana at $4; and Zoe value each apple at $2 and each banana at $1. Consider the following three allocations Allocation 1: Xavier has 4 apples and 3 bananas, Yolanda has 2 apples and 2 bananas, and Zoe has 3 bananas and 4 apples. Allocation 2: Xavier has 3 apples and 4 bananas, Yolanda has 2 apples and 2 bananas, and Zoe has 4 bananas and 3 apples. Allocation 3: Xavier and Zoe have nothing, Yolanda has 9 apples and 9 bananas. Rank the three allocations according to the total surplus criterion Allocation 3 is the best, and Allocation 2 is better than Allocation 1. Allocation 3 is the worst, Allocations 1 and 2 are tied. Allocation 1 is the best, Allocations 2 and 3 are tied. It is impossible to compare any two allocations. Allocation 3 is the best, Allocations 1 and 2 are tied.
Answer #1

Allocation 1 surplus:

Xavier surplus=(4 apple * 1$) + (3 bananas * 2$). = 4 + 6 = 10$

Yolanda surplus = ( 2 apples * 2$) + (2 bananas * 4$) = 4+8

= 12$

Zoe surplus = (3 bananas *1$)+(4 apples*2$) = 3+8 = 11$

Total surplus = 10+12+11= 33$.

Allocation 2 surplus:

Xavier surplus=(3 apple * 1$) + (4 bananas * 2$). = 3 + 8 = 11$

Yolanda surplus = ( 2 apples * 2$) + (2 bananas * 4$) = 4+8

= 12$

Zoe surplus = (4 bananas *1$)+(3 apples*2$) = 4+6= 10$

Total surplus =11+12+10 = 33$

Allocation 3 surplus:

Both Xavier and Zoe has 0 surplus.

Yolanda surplus = (9 bananas *4$)+(9 apples*2$) =36+18 = 54$

Total surplus =0+0+54 = 54$.

Hence allocation 3 is the best, Allocation 1 and 2 are tied. This is because in the above calculation, allocation 1 surplus and allocation 2 surplus are same even though each surplus of both Zoe and Xavier are different in allocation 1 and 2. In allocation 3, even though both Xavier and Zoe has 0 surplus, surplus of Yolanda makes the total surplus of allocation 3 more than the total allocation surplus of both 1 and 2.